What is the WACC of a company with the following capitalization. The company is in the 33% tax bracket, and floatation costs add 1.2% from the firm's perspective.

Allocation Market Cost/ Expected Return
Debt 0.3 5.5%
Retained Earnings 0.4 9.4%
Common Stock (new issue)
Answer #1

Gross cost of debt = 5.5%

Net cost of debt = (1 - tax rate)x Gross cost of debt = (1 - 0.33)x 5.5% = 3.685% and  alllocation 0.3

Cost of retained earnings = 9.4% and allocaton = 0.4

Cost of new issue = 9.4 + 1.2 = 10.6% and allocation = 0.3

Therefore WACC = 0.3x3.685 + 0.4x 9.4 + 0.3x10.6 = 8.0455%